OpenStudy (anonymous):

Substitution (x + 2y = 10) (x - y = 4)

OpenStudy (anonymous):

x+2y=10 x-y=4 x=10-2y (10-2y)-y=4 10-3y=4 -3y=-6 y=2 x=6

OpenStudy (anonymous):

Ykz is correct